# RELATE: Relation Extraction in Biomedical Abstracts with LLMs and Ontology Constraints

### v1

@@ 4 Experiments
tab:chemprot_overall LLM Embedding Exact Match a@1 a@3 a@5 a@10 MRR MedGemma 27B nomic 0.464 0.528 0.686 0.744 0.755 0.612 MedGemma 27B bge-m3 0.467 0.468 0.686 0.788 0.812 0.590 MedGemma 27B nomic + SapBERT 0.520 0.100 0.328 0.688 0.914 0.300 MedGemma 27B bge + SapBERT 0.516 0.100 0.321 0.691 0.940 0.302
Table 1: RELATE performance on ChemProt. Exact Match: accuracy of RELATE; a@$k$: accuracy in top-$k$ candidates before LLM-based refinement in Stage 3 of RELATE; MRR: Mean reciprocal rank.
We assess RELATE on two independent datasets with different ontology models, as well as four different embedding model configurations.
4.1 Datasets
The ChemProt dataset (Krallinger et al., 2017) provides annotated chemical-protein relations representing diverse pharmacological and biochemical relationship types, including regulators, modulators, agonists, antagonists, and substrate interactions. We treat the information in ChemProt as ground truth abstract-predicate pairs for evaluation purposes. Standard metrics in information retrieval are used, such as exact match accuracy, accuracy@$k$, and Mean Reciprocal Rank (MRR). Recall from Figure LABEL:fig:reranking_prompt that Stage 3 of RELATE only outputs its top choice predicate $r^{*}$ or NONE, and does not return a fully re-ranked or re-ordered list of $k$ candidates. Therefore, exact match accuracy is calculated with respect to the final output of RELATE, $r^{*}$, but accuracy@$k$ and MRR are calculated using the top-$k$ candidates after Stage 2, before LLM-based Stage 3. This will allow us to analyze the effect of LLM-based refinement in isolation as well.
Next, we apply RELATE to 2,400 PubMed abstracts from the HEAL Project [^4] https://heal.nih.gov targeting opioid use disorder research. This real-world dataset reflects the practical application scenario where relationships extracted from literature must be mapped to standardized Biolink predicates for integration into a large biomedical KG [^5] Name hidden during double blind review process.. Unlike ChemProt, the HEAL abstracts lack ground-truth annotations, requiring qualitative assessment of RELATE’s outputs.
Importantly, this corpus captures the complexity and noisiness of real biomedical literature. In contrast to benchmark datasets, these abstracts often contain methodological statements or negative scientific findings (e.g., drug does not affect condition, gene not associated with disease). Such cases underscore the necessity of RELATE’s rejection and negation-handling capabilities to ensure only semantically appropriate biomedical relationships enter the KG. As part of the integration plan, the abstract authors and Subject Matter Experts (SMEs) are being recruited to systematically evaluate the quality and reliability of the extracted triples.
4.2 LLM and Embedding Models
We experimented with MedGemma (4B), LLaMA2 (7B), LLaMA3 (8B), LLaMA3 (70B), MedGemma (27B), and OpenAI GPT-4o-mini (unknown number of parameters). The last two demonstrated greater consistency and fewer hallucinations compared to the other models. Since those two models had very similar performances, we only report results using MedGemma in this work. However, note that any other LLM could be substituted to reproduce the pipeline. MedGemma-27B is Google’s medical language model (Sellergren et al., 2025), trained exclusively on medical text and optimized for efficient inference through architectural improvements in performance and computational efficiency.
As with the LLMs, we also experimented with several embedding approaches. While we prioritize accessibility through ollama-based open-access models, we tested OpenAI’s text-embed-large models and others with 8,192-token context windows that provide general-purpose semantic understanding. These include nomic-embed-text (Zach et al., ), a 768-dimensional embedding model, and bge-m3, a 1,024-dimensional multilingual embedding model for diverse textual contexts (Chen et al., 2024b). Since nomic-embed-text and bge-m3 achieved comparable performance, we report only the results obtained with the 768-dimensional nomic-embed-text in this work.
For (optional) hybrid retrieval, we experimented with adding fine-tuned SapBERT (Liu et al., 2021). We initialized the model with BiomedNLP-PubMedBERT-base-uncased-abstract -fulltext and optimized it using a contrastive learning framework. Specifically, we employed multi-similarity loss with hard negative mining, which encourages the model to better distinguish between semantically close but incorrect pairs. To keep training efficient while still capturing relevant semantic features, we limited input sequences to 25 tokens and represented each instance using the [CLS] embedding.
@@ 5 Results and Discussions
These results highlight a trade-off: baseline embeddings rank the correct answer more accurately at the top, while SapBERT embeddings broaden coverage and guarantee inclusion within candidate sets. RELATE resolves this through its LLM-based refinement stage, which reranks SapBERT’s expanded candidate lists in context. This allows RELATE to recover early precision while preserving high recall, transforming broad retrieval into precise ontology-constrained mappings. In short, SapBERT enhances retrieval coverage, but RELATE’s LLM-based refinement is essential to convert this coverage into accurate, ontology-aligned relations.
@@
RELATE’s rejection capability (NONE) represents a critical quality control mechanism. Out of 2,400 PubMed abstracts from the HEAL project, 10 relationships were appropriately rejected as unsuitable for Biolink predicate mapping (rejection rate:  0.4%). Analysis of these rejections reveals contextual understanding that distinguishes between legitimate biomedical relationships and other types of associations commonly found in scientific literature.
@@
These rejections occurred despite high embedding similarity scores, demonstrating the reranker’s comprehensive contextual analysis. The low rejection rate (0.4%) indicates that the majority of extracted relationships represent valid biomedical associations while maintaining strict quality standards for edge cases requiring specialized handling.
Chemprot analysis also reveals potential false rejections where high similarity scores failed to produce mappings. Cases like “loperamide modifies corticotropin-releasing hormone actio“ (0.661 score for “modulator“) and “ammonium sulfate reduces CBG concentrations“ (0.632 for “downregulator“) were rejected despite appropriate candidates. While some discrepancies reflect stricter semantic criteria, others suggest overly conservative contextual reranking, highlighting evaluation challenges with traditional benchmarks.
@@
Out of 2,400 PubMed abstracts from the HEAL project, 77 relationships were correctly flagged as negated assertions, revealing how well the system processes the nuanced ways of reporting biomedical negative findings. Most prevalent was “not associated with”, which accounted for nearly half of the negated cases (32 instances). Beyond these straightforward cases, we found the pipeline successfully navigating more complex terrain. Other negations like “does not affect” appeared in 8 cases, while direct causal negations (“does not cause”) showed up 7 times. The remaining quarter of cases presented genuine linguistic challenges. Consider phrases like “remained impaired despite treatment” or “not necessary for acute insulin-induced uptake”. These require contextual understanding that goes well beyond pattern matching. The system had to recognize that “fails to affect” and “unable to deliver” carry negative semantic weight, even when the word “not” never appears.
@@
To evaluate SapBERT’s effect, we compared 2,139 predicate assignments between bge-m3 and bge-m3 + SapBERT. Their disagreement patterns reveal four key improvements by SAPBERT aligned with biomedical expert judgment. Clinical terminology refinement was most prominent, with 28 cases shifting from associated with: increased likelihood of to predisposes to condition. Domain specificity enhancement appeared in 12 cases where generic associated with became gene associated with condition for relationships like “DRD4 associated with Heroin Addiction,” recognizing genetic contexts require more informative predicates.
@@ 6 Conclusion
We introduced RELATE, a three-stage pipeline for automated ontology-constrained predicate mapping in biomedical relation extraction. By combining ontology-driven preprocessing, efficient vector search, and context-aware LLM reranking, RELATE systematically converts free-text relations into ontology-aligned, interoperable knowledge graph edges. Experiments on ChemProt demonstrate the trade-offs between early precision and retrieval depth, while application to 2,400 HEAL Project abstracts illustrates RELATE’s ability to handle noisy real-world biomedical literature, including identification of negated assertions and appropriate rejection of non-biomedical associations. Together, these results establish a foundation for scalable, ontology-constrained relation extraction and highlight a path forward for enriching biomedical knowledge graphs with high-quality, standardized information. Future work includes improving computational overhead, designing a scalable evaluation that combines expert judgment with automated consistency checks, and integration into the existing large biomedical KG (ROBOKOP).
@@
We thank the LitCoin project partners for their immense contributions, the HEAL initiative consortium for providing access to opioid research abstracts. Also, we acknowledge the RENCI’s Data Analytics group and Knowledge Graph Working Group for constructive reviews. This research was supported by NIH #75N95023C00032 (LITCOIN)
Appendix A Relation Extraction Pipeline Stages Algorithms
Input: Biolink Model YAML $\mathcal{B}$, Biolink Toolkit $\mathcal{T}$, LLM $\mathcal{M}$, Embedding model $f_{\theta}$
Output: Embedding store $\mathcal{V}$, Positive descriptors $\mathcal{D}^{+}$, Negative descriptors $\mathcal{D}^{-}$
@@ Appendix A Relation Extraction Pipeline Stages Algorithms
for each predicate mapping $r\in\mathcal{R}$ do
 $\text{predicate}\leftarrow r.\text{predicate}$;
 $\text{descriptions}\leftarrow$ GatherDescriptions(predicate)($\text${predicate});
 $\mathcal{D}^{+}[\text{predicate}]\leftarrow\text{descriptions}$;
@@
 for each description $d\in\text{pos\_descriptions}$ do
@@
    $d^{neg}\leftarrow\mathcal{M}(\text{prompt})$;
    $\text{neg\_descriptions}\leftarrow\text{neg\_descriptions}\cup\{d^{neg}\}$;
@@
 $\text{neg\_key}\leftarrow p+\text{"\_NEG"}$;
@@
 $\mathbf{v}_{p}\leftarrow f_{\theta}(\text{combined\_text})$;
 $\mathcal{V}\leftarrow\mathcal{V}\cup\{(\text{label}:p,\text{embedding}:\mathbf{v}_{p})\}$;
@@
return $\mathcal{V},\mathcal{D}^{+},\mathcal{D}^{-}$;
@@
Input: Loose triples $\mathcal{T}=\{(s_{i},o_{i},r_{i},a_{i})\}_{i=1}^{n}$, Predicate embeddings $\mathcal{V}$, Embedding model $f_{\theta}$, Top-k parameter $k$
Output: $\mathcal{C}$
@@
$\mathcal{R}_{embed}\leftarrow\emptyset$;
for each triple $(s_{i},o_{i},r_{i},a_{i})\in\mathcal{T}$ do
 $\mathbf{v}_{r_{i}}\leftarrow f_{\theta}(r_{i})$;
 $\mathcal{R}_{embed}\leftarrow\mathcal{R}_{embed}\cup\{(i,\mathbf{v}_{r_{i}})\}$;
@@
for each $(i,\mathbf{v}_{r_{i}})\in\mathcal{R}_{embed}$ do
@@
 for each $(\text{label}_{j},\mathbf{v}_{j})\in\mathcal{V}$ do
    $\text{sim}_{j}\leftarrow\cos(\mathbf{v}_{r_{i}},\mathbf{v}_{j})=\frac{\mathbf{v}_{r_{i}}\cdot\mathbf{v}_{j}}{\|\mathbf{v}_{r_{i}}\|\|\mathbf{v}_{j}\|}$;
    $\text{similarities}\leftarrow\text{similarities}\cup\{(\text{label}_{j},\text{sim}_{j})\}$;
@@
 $\text{sorted}\leftarrow$ SortDescending(similarities)($\text${similarities});
 $\text{candidates}_{i}\leftarrow\text{sorted}[1:k]$;
 $\mathcal{C}\leftarrow\mathcal{C}\cup\{(i,\text{candidates}_{i})\}$;
@@
return $\mathcal{C}$;
@@
Input: Loose triples $\mathcal{T}=\{(s_{i},o_{i},r_{i},a_{i})\}_{i=1}^{n}$, Candidate sets $\mathcal{C}$, LLM $\mathcal{M}$
Output: Mapped triples $\mathcal{T}^{*}=\{(s_{i},o_{i},r_{i}^{*},a_{i})\}_{i=1}^{n}$
$\mathcal{T}^{*}\leftarrow\emptyset$;
for each triple $(s_{i},o_{i},r_{i},a_{i})\in\mathcal{T}$ do
 $\text{candidates}_{i}\leftarrow\mathcal{C}[i]$;
@@
 // Prompt Construction
 $\text{prompt}\leftarrow$ ”Given the biomedical relationship context:”;
 $\text{prompt}\leftarrow\text{prompt}+\text{" Subject: "}+s_{i}+\text{" Object: "}+o_{i}$;
 $\text{prompt}\leftarrow\text{prompt}+\text{" Relationship: "}+r_{i}+\text{" Abstract: "}+a_{i}$;
 $\text{prompt}\leftarrow\text{prompt}+\text{" Select best Biolink predicate or NONE:"}$;
 for $j=1$ to $|\text{predicate\_list}|$ do
    $\text{prompt}\leftarrow\text{prompt}+j+\text{". "}+\text{predicate\_list}[j]$;
   end for
 $\text{prompt}\leftarrow\text{prompt}+(|\text{predicate\_list}|+1)+\text{". NONE"}$;
@@
 $\text{response}\leftarrow\mathcal{M}(\text{prompt})$;
[…]
 if $r_{i}^{*}=\text{"NONE"}$ or $r_{i}^{*}\notin\text{predicate\_list}$ then
@@
 $\mathcal{T}^{*}\leftarrow\mathcal{T}^{*}\cup\{(s_{i},o_{i},r_{i}^{*},a_{i})\}$;
@@
return $\mathcal{T}^{*}$;

### v2

@@ 4.2 LLM and Embedding Models
tab:chemprot_overall Embedding Model Stage 2 (Candidate Retrieval) Post-Refinement a@1 a@3 a@5 a@10 MRR Exact Match (a@1) nomic 0.528 0.686 0.744 0.755 0.612 0.464 bge-m3 0.468 0.686 0.788 0.812 0.590 0.467 nomic + SapBERT 0.100 0.328 0.688 0.914 0.300 0.520 bge-m3 + SapBERT 0.100 0.321 0.691 0.940 0.302 0.516 Baseline (no embedding) - - - - - 0.220
Table 1: RELATE performance on ChemProt. Stage 2 (Candidate Retrieval): a@$k$ = accuracy in top-$k$ candidates before LLM-based refinement; MRR = mean reciprocal rank. Post-Refinement: Exact Match = final accuracy after LLM-based refinement; Baseline = direct LLM classification without RELATE’s retrieval-refinement pipeline. LLM used is MedGemma (27B) for all.
@@ 5 Results and Discussions
RELATE Pipeline is Necessary.
To validate RELATE’s pipeline design, we compared against direct LLM classification where MedGemma (27B) selects from the 9 ChemProt predicates without retrieval-refinement. The baseline achieves only 22.0% exact match (Table LABEL:tab:chemprot_overall), compared to RELATE’s 52.0% (136% improvement). This demonstrates that structured retrieval outperforms asking an LLM to select directly from a predicate list. The baseline’s poor performance reflects the difficulty of simultaneous predicate understanding and selection without semantic guidance from embedding-based retrieval.
The exact match accuracy of 52% reflects the difficulty of mapping to fine-grained ontology predicates. However, this metric may understate practical utility. Manual analysis of the misclassifications show semantically acceptable near-misses, such as mapping to agonist instead of modulator. When allowing such semantically close alternatives, accuracy increases to approximately 70-80%.
@@
These results highlight a trade-off: baseline embeddings rank the correct answer more accurately at the top, while SapBERT embeddings broaden coverage and guarantee inclusion within candidate sets.
While selecting the top-ranked candidate from embeddings without LLM refinement would yield higher exact match on this benchmark, such an approach assumes embedding similarity reliably ranks correct predicates first across all datasets and relation types. This assumption cannot be guaranteed a priori. RELATE’s design prioritizes robustness through SapBERT’s comprehensive retrieval combined with LLM-based contextual reranking, which serves as a buffer against cases where embedding similarity alone performs poorly. This allows RELATE to recover early precision while preserving high recall, transforming broad retrieval into precise ontology-constrained mappings regardless of how initial embeddings behave on new corpora.
@@
RELATE’s rejection capability (NONE) represents a critical quality control mechanism. Out of 2,400 PubMed abstracts from the HEAL project, 10 relationships were appropriately rejected as unsuitable for Biolink predicate mapping (rejection rate:  0.4%). Manual analysis of these rejections reveals contextual understanding that distinguishes between legitimate biomedical relationships and other types of associations commonly found in scientific literature.
@@
These rejections occurred despite high embedding similarity scores, demonstrating the reranker’s comprehensive contextual analysis. The low rejection rate (0.4%) suggests that the majority of extracted relationships represent valid biomedical associations while maintaining strict quality standards for edge cases requiring specialized handling. All 10 rejections were deemed appropriate (100% precision).
On ChemProt, there were 5 correct rejections out of 18 total rejections. These false rejections occur where, despite high similarity candidate scores, RELATE failed to produce mappings. Cases like “loperamide modifies corticotropin-releasing hormone action” (0.661 score for modulator) and “ammonium sulfate reduces CBG concentrations” (0.632 for downregulator) were rejected despite appropriate candidates. While some discrepancies reflect stricter semantic criteria, others suggest overly conservative contextual reranking, highlighting evaluation challenges with traditional benchmarks. These contrasting patterns reflect differences in semantic alignment between upstream extraction and ontology predicates across datasets. See more details in Appendix C.2.
@@
Out of 2,400 PubMed abstracts from the HEAL project, 77 relationships were correctly flagged as negated assertions, revealing how well the system processes the nuanced ways of reporting biomedical negative findings. Most prevalent was “not associated with”, which accounted for nearly half of the negated cases (32 instances). Beyond these straightforward cases, we found the pipeline successfully navigating more complex terrain. Other negations like “does not affect” appeared in 8 cases, while direct causal negations (“does not cause”) showed up 7 times. The remaining quarter of cases presented genuine linguistic challenges. Consider phrases like “remained impaired despite treatment” or “not necessary for acute insulin-induced uptake.” These require contextual understanding that goes well beyond pattern matching. The system had to recognize that “fails to affect” and “unable to deliver” carry negative semantic weight, even when the word “not” never appears. See more details in Appendix C.3.
@@
To evaluate SapBERT’s effect, we compared 2,139 predicate assignments between bge-m3 and bge-m3 + SapBERT. Their disagreement patterns reveal four key improvements by SAPBERT aligned with biomedical expert judgment. Clinical terminology refinement was most prominent, with 28 cases shifting from associated with: increased likelihood of to predisposes to condition. Domain specificity enhancement appeared in 12 cases where generic associated with became gene associated with condition for relationships like “DRD4 associated with Heroin Addiction,” recognizing that genetic contexts require more informative predicates.
@@
Limitations.
RELATE’s performance depends on the quality of upstream entity and relation extraction. While we do not assume perfect extraction, the input quadruples must be reasonably accurate. Errors in entity boundaries, entity type misclassifications, or poorly extracted relation phrases will propagate through the standardization pipeline. In production KG construction, additional methods address these entity errors. RELATE’s modular design allows it to benefit from improvements in upstream extraction systems without requiring system redesign.
@@ 6 Conclusion
We introduced RELATE, a three-stage pipeline for automated ontology-constrained predicate mapping in biomedical relation extraction. By combining ontology-driven preprocessing, efficient vector search, and context-aware LLM-based reranking, RELATE systematically converts free-text relations into ontology-aligned and interoperable KG edges. Experiments on ChemProt demonstrate the trade-offs between early precision and retrieval depth, while application to 2,400 HEAL Project abstracts illustrates RELATE’s ability to handle noisy real-world biomedical literature, including detecting negated assertions and appropriately rejecting non-biomedical associations. Together, these results establish a foundation for scalable, ontology-constrained relation extraction and highlight a path forward for enriching biomedical KGs with high-quality, standardized information. Future work includes improving computational overhead, incorporating further top-$k$ reordering and evaluation in the refinement stage, designing a scalable evaluation that combines expert judgment with automated consistency checks, and integration into the existing large biomedical KG (ROBOKOP). As part of the integration plan, the HEAL abstract authors and domain experts are being recruited to systematically evaluate the quality and reliability of RELATE and the LLM-extracted quadruples that serve as its inputs.
@@
We thank the LitCoin project partners for their immense contributions, and the HEAL initiative consortium for providing access to opioid-related research abstracts. Also, we acknowledge the RENCI’s Data Analytics group and Knowledge Graph Working Group for constructive reviews. This research was supported by NIH #75N95023C00032 (LITCOIN).
Appendix A Clinical and Research Applications
A.1 RELATE’s Role and Significance
RELATE is a relation standardization component that ensures relations extracted from biomedical text are accurately mapped to standardized ontology predicates. This standardization is foundational for downstream knowledge graph applications.
In literature-based discovery, standardized predicates enable systematic querying across diverse sources. A researcher investigating drug repurposing can query for all compounds that upregulate proteins downregulated in a specific disease, confident that the upregulator predicate maintains consistent semantics across thousands of papers. Without standardization, the same relationship might appear as “increases expression,” “upregulates,” or “enhances activity,” requiring separate queries and potentially missing relevant findings.
Biomedical knowledge graph systems such as ROBOKOP require ontology-aligned relations for reliable reasoning and inference. RELATE ensures that literature-derived evidence integrates seamlessly with structured knowledge bases, enabling complex queries that traverse multiple relationship types and data sources.
Our evaluation on HEAL demonstrates how relations from distinct domains can be mapped to common Biolink predicates, while ChemProt enables unified knowledge graphs that integrate chemical interactions.
A.2 Comparison with End-to-End Systems
Unlike end-to-end systems such as AutoRD, which perform extraction, standardization, and graph construction, RELATE focuses specifically on standardization. This provides modularity, allowing RELATE to work with any upstream extraction system, and scalability, handling 284 Biolink predicates compared to 6 relation types in AutoRD.
A.3 Broader Applicability
While evaluated on biomedical corpora, RELATE generalizes to any domains with structured predicate vocabularies for computing embeddings.
Appendix B RELATE Stages Algorithms
Algorithms 1, 2, and 3 summarize the three stages of RELATE. When the optional SAPBERT workflow enhancement is on, RELATE follows the modified pipeline described in Figure LABEL:fig:workflowplus.
Input: Biolink Model YAML $\mathcal{B}$, Biolink Toolkit $\mathcal{T}$, Embedding model $f_{\theta}$, LLM
@@ Appendix B RELATE Stages Algorithms
for each predicate $r\in\mathcal{R}$ do
 $\text{descriptions}\leftarrow$ GatherDescriptions(r)(r);
 $\mathcal{D}^{+}[r]\leftarrow\text{descriptions}$;
@@
 for each description $d\in\mathcal{D}^{+}[p]$ do
@@
    $d^{-}\leftarrow LLM(\text{prompt})$;
    $\text{neg\_descriptions}\leftarrow\text{neg\_descriptions}\cup\{d^{-}\}$;
@@
 $\text{neg\_key}\leftarrow p+\text{``\_NEG"}$;
@@
 $v\leftarrow f_{\theta}(\text{combined\_text})$;
 $\mathcal{V}\leftarrow\mathcal{V}\cup\{(\text{label}:p,\text{embedding}:v)\}$;
@@
return Embedding store $\mathcal{V}$, Positive descriptors $\mathcal{D}^{+}$, Negative descriptors $\mathcal{D}^{-}$;
@@
Input: Loose quadruples $\{(s_{i},o_{i},T_{i},a_{i})\}$, Predicate embeddings $\mathcal{V}$, Embedding model $f_{\theta}$, Parameter $k$
@@
$\mathcal{Q}\leftarrow\emptyset$;
for each quadruples $(s_{i},o_{i},T_{i},a_{i})$ do
 $q_{i}\leftarrow f_{\theta}(T_{i})$;
 $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(i,q_{i})\}$;
@@
for each $(i,q_{i})\in\mathcal{Q}$ do
@@
 for each $(p,v)\in\mathcal{V}$ do
    $\text{sim}\leftarrow\cos(q_{i},v)$;
    $\text{similarities}\leftarrow\text{similarities}\cup\{(p,\text{sim})\}$;
@@
 $\mathcal{C}[i]\leftarrow$ Top-kk(similarities)($\text${similarities});
@@
return Candidate sets $\mathcal{C}$;
@@
Input: Loose quadruples $\{(s_{i},o_{i},T_{i},a_{i})\}$, Candidate sets $\mathcal{C}$, LLM
$\mathcal{A}\leftarrow\emptyset$;
for each quadruple $(s_{i},o_{i},T_{i},a_{i})$ do
@@
[…]
@@
 $r_{i}^{*}\leftarrow LLM(\text{prompt})$;
 if $r_{i}^{*}=\texttt{NONE}$ or $r_{i}^{*}\notin\text{predicate\_list}$ then
@@
 $\mathcal{A}\leftarrow\mathcal{A}\cup\{(s_{i},o_{i},r_{i}^{*},a_{i})\}$;
@@
return Mapped quadruples $\mathcal{A}=\{(s_{i},o_{i},r_{i}^{*},a_{i})\}$;